What will be resistivity of a metal wire of 2m length and 0.6 mm in diameter, if the resistance of the wire is 50




1
Expert's answer
2021-06-28T17:09:03-0400

l= 2 m

d = 0.6 mm

R= 50 Ω

Radius of wire "r = 0.3\; mm = 0.3 \times 10^{-3} \;m"

Area of cross section:

"A = πr^2 \\

A = 3.14 \times (0.3 \times 10^{-3})^2 = 0.2826 \times 10^{-6}\;m^2 \\

R = ρ \times \frac{l}{A} \\

50 = ρ \times \frac{2}{0.2826 \times 10^{-6}} \\

ρ = \frac{50 \times 0.2826 \times 10^{-6}}{2} = 7.065 \times 10^{-6} \;Ωm"